OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. 1. (1 pt) Which related equation can be used to solve the equation? Column A Column B 1. m + 5 = –7.3 2. m + 5 = 7.3 3. m + 7.3 = –5 4. m + 7.3 = 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

your columns didn't really format very well, kinda hard to read

OpenStudy (misty1212):

\[m + 5 = –7.3\iff m=-7.3-5\] i guess that is what you are looking for

OpenStudy (anonymous):

1 2 3 4 in column a

OpenStudy (texaschic101):

I just did these in your last question

OpenStudy (anonymous):

its different

OpenStudy (texaschic101):

1. m + 5 = -7.3 m = -7.3 - 5 2. m + 5 = 7.3 m = 7.3 - 5 3. m + 7.3 = -5 m = -5 - 7.3 4. m + 7.3 = 5 m = 5 - 7.3

OpenStudy (anonymous):

thank you